OpenStudy (anonymous):

PLEASE HELP! Use a double-angle formula to write the given expression as a single trigonometric function of twice the angle. 1 − 2 sin^2 π/3

OpenStudy (anonymous):

that is the formula for \(\cos(2\theta)\) so \[\cos(\frac{2\pi}{3})\]